OpenStudy (anonymous):

What is the vertex of the graph of y = −2x2 + x + 3?

OpenStudy (anonymous):

a. ordered pair 1 over 4 comma 25 over 8 b. ordered pair negative 1 over 4 comma 27 over 8 c. ordered pair 1 over 4 comma 27 over 8 d. ordered pair negative 1 over 4 comma 25 over 8

Mehek (mehek14):

use the formula \(\tt{x=\dfrac{-b}{2a}}\)

OpenStudy (anonymous):

-1/-4?

Mehek (mehek14):

so x = 1/4

Mehek (mehek14):

now put 1/4 for all of the x's into the equation to find y

OpenStudy (anonymous):

.25 + 1/4 + 3

OpenStudy (anonymous):

3.5?

Mehek (mehek14):

3.125

Mehek (mehek14):

which in fraction form is 25/8

OpenStudy (anonymous):

wouldnt it be (1/4, 27/8)

Mehek (mehek14):

no 27/8 = 3.375 not 3.125

OpenStudy (anonymous):

or 25/8?

Mehek (mehek14):

25/8

OpenStudy (anonymous):

thank you!

Mehek (mehek14):

yw :)